How long must one wait for $1250 to grow to $3000 at 12% compounded monthly?
Answer #1

12% compounded monthly becomes (12/12)% = 1% effective rate monthly.

Thw formula for future value is given as

Future value = Present value *(1+interest rate)^number of periods

Thus, 3000 = 1250*(1+0.01)^number of months

So, ln(3000/1250) = number of months*ln(1.01)

Numbef of months = 87.9839

Thus, 88 months are required for $1250 to grow to $3000 at the given rate.

Comment in case of any queries.